how old are you daddy said the son father replied 12 year hence you will be as old as i was 12years ago andi shall be 3 times as old as you were 12 years ago find their present ages

Answer by ankor@dixie-net.com(22740)

"How old are you daddy?" said the son.
Father replied, "12 year hence, you will be as old as I was 12 years ago, and
I shall be 3 times as old as you were, 12 years ago."
Find their present ages.
:
It makes a lot more sense if you use some punctuation here.
:
Let d = Dad's age
Let s = Son's age
:
"12 year hence, you will be as old as I was 12 years ago,"
s + 12 = d - 12
s = d - 12 - 12
s = d - 24
:
"and I shall be 3 times as old as you were, 12 years ago."
d + 12 = 3(s-12)
d + 12 = 3s - 36
d = 3s - 36 - 12
d = 3s - 48
replace s with (d-24) from the 1st equation
d = 3(d-24) - 48
d = 3d - 72 - 48
d = 3d - 120
120 = 3d - d
120 = 2d
d = 120/2
d = 60 yrs is Dad's present age
and
s = d - 24
s = 60 - 24
s = 36 yrs is Son's present age
:
:
You can check this out in the given statements.